OpenStudy (anonymous):

Which function has the greatest rate of change? A. 14x – 2y = 8 B. y = 6x – 5 C. x y −2 −7 −1 −2 0 3 D. A linear function that goes through point 1, 3 and point 0, negative 2

OpenStudy (anonymous):

Please Help! ^.^

OpenStudy (butterflydreamer):

rate of change means slope :) So you want to find the slope for each of the functions.

OpenStudy (anonymous):

Okay! Thank You! ⌒‐⌒

OpenStudy (anonymous):

So what would the answer be?

OpenStudy (butterflydreamer):

you're welcome :) Well can you tell me what you got as the slope for each of the functions?

OpenStudy (welshfella):

rearrange the functions to the form y = mx + b m will be your slope

OpenStudy (welshfella):

B is in that form already y = 6x - 5 so its slope = 6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

how do we solve a,c and d?

OpenStudy (welshfella):

well rearrange A to y = mx + b form

OpenStudy (anonymous):

Slope = 7

OpenStudy (welshfella):

yes now for C you need to see how y increases with unit increase of x as x goes from -2 to -1 y does from -7 to -2 so thats a slope of ?

OpenStudy (butterflydreamer):

you might want to use the slope formula: m (slope) = rise/ run \[m = \frac{ y2 - y1 }{ x2 - x1 }\]

OpenStudy (welshfella):

yes that will give you the slope from the points on the graph

OpenStudy (welshfella):

you can use that for option D

OpenStudy (welshfella):

for option C the slope is -2-(-7) / -1-(-2)

OpenStudy (anonymous):

-7 for C?

OpenStudy (welshfella):

no

OpenStudy (welshfella):

how mant units is it from -7 up to -2 ?

OpenStudy (anonymous):

5

OpenStudy (welshfella):

right and -2 to -1 is one unit slope = 5/1

OpenStudy (welshfella):

now to find the slope for D plug in the points (1,3) and (0,-2) into (x2,y2) and (x1,y1) into the formula for m that butterfly dreamer posted

OpenStudy (anonymous):

Thank You! :)

OpenStudy (welshfella):

yw